OpenStudy (anonymous):

the edges of a tetrahedron measure 3,4,5,6,7 and 8 units. find the edges of a similar tetrahedron containing 27 times the volume. HEELLP ME PLLEEAASSEE!

Directrix (directrix):

If two solids are similar, the cube of the scale factor of the two solids is equal to the ratio of the volumes. Let x be the length of the side of the larger, similar solid. Let be the side that corresponds to 3 in the smaller solid. According to the theorem: (1/x)^3 = (1/27) (1/x)^3 = (1/3)^3 So, (1/x) = (1/3) The scale factor from small to large is one to three. So, triple the given edges of the smaller solid to get the lengths of the larger solid. @sassy016